# Evaluating the Utility of Conformal Prediction Sets for AI-Advised Image Labeling

### v1

@@ 4.4. Willingness-to-Pay
We present the distributions of elicited willingness-to-pay (WTP) measures in Figure S2 with summary statistics in Table S4 of Appendix B.2. Elicited WTPs across conditions are quite similar, suggesting the measure may be noisy. Participants are willing to pay slightly more for Top-10 (; CI: ) than RAPS (; CI: ) and Top-1 (; CI: ).
Nonetheless, as detailed in Section 3.6, we quantify participants’ perceived utility for each prediction display using a measure we devised, called “willingness-to-overpay”. This proxy focuses on economic valuation, offering direct monetary comparisons between the monetary value of a prediction display that participants perceive (i.e., willingness-to-pay) against how much additional expected bonus participants gained from having access to each prediction display.
| Condition | Accuracy | Expected Bonus | Expected Bonus Difference |
| baseline | 0.41 | 1.64 | 0.00 |
| Top-1 | 0.63 | 2.52 | 0.86 |
| Top-10 | 0.64 | 2.56 | 0.88 |
@@
Bootstrapped 95% confidence interval for willingness-to-overpay
We note that this measure is imperfect as a comparator for any individual willingness-to-pay, because we cannot observe the counterfactual performance without a prediction display for each participant individually. In aggregate, however, Figure 8 suggests that participants may over-value the prediction displays to a similar extent across conditions. If anything, we see weak evidence that RAPS participants, despite benefiting more from using the prediction sets, perceive a lower monetary value from the prediction set, relative to Top- predictions. RAPS participants are willing to overpay by on average (CI ), slightly less than Top-1 participants (, CI ). Participants find Top-10 prediction most useful on average (; CI ).
@@ 4.5. Qualitative Analysis of Strategies
From our open coding, we find that participants employed different strategies between prediction displays (see Figure S3 in Appendix B.3). In general, most participants reported consulting predictions when decision-making ( out of 450). Some participants reported first inspecting predictions and assessing their accuracy using their judgment (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150), while others report relying first on their own intuitions to form a general idea, and then seeing if it aligned with any predictions (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150). Very few participants chose to rely completely on their own intuitions because they find predictions inaccurate or not useful (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150). Others reported they would only consult predictions when the image was too hard to recognize and they had no clue (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150). Among participants who reported that they trusted AI prediction(s), some of them in Top- conditions placed higher trust on predictions with high softmax pseudo-probability (e.g., ) (Top-1: out of 23; Top-10: out of 35), while those in RAPS condition focused more on the top suggestions in the prediction set ( out of 26).
Collectively participants used all features of our interface, reporting that the dropdown ( out of 599), bottom-up ( out of 599), and keyword search ( out of 599) methods were useful in helping them identify a label, and the representative images were informative for aiding understanding of the meaning of labels ( out of 599). Of participants who viewed prediction displays, those in the Top-1 condition appeared most likely to use the label space hierarchy network to find a better match ( out of 150, compared to out of 150 and out of 150).
@@ 5. Discussion
Conformal prediction has achieved recent visibility as a distribution-free approach to quantifying uncertainty in model predictions, which is robust to some forms of model misspecification. Our results suggest that decision-makers can use prediction sets effectively in an AI-advised labeling task, but that their benefits over status quo Top- presentation displays depend on the properties of the setting. In real-world scenarios where the deployed model is typically well calibrated and the predictions are reliable, we find the size of the prediction set is a critical determinant of its utility. Uncertainty quantification by prediction sets appears to be more useful with a smaller set size and less effective with a larger one that can increase users’ cognitive load. In contrast, when the model encounters unforeseen and challenging OOD instances that can cause Top- predictions and their associated uncertainty scores unreliable, prediction sets have a comparative advantage, regardless of their set size, at least in a setting like ours where coverage guarantees are maintained. We elucidate our main findings and discuss the broader implications of our study.
@@
Our results suggest that the utility of conformal prediction sets is linked to set size, presumably due to an underlying factor of cognitive load. A large set embodying more uncertainty requires users to navigate and evaluate many incorrect predictions before finding the correct answer. This increase in cognitive load makes prediction sets less effective and may encourage more satisficing (Simon, 1956), resulting in low response accuracy. Our results show that labeling accuracy using prediction sets is higher than those of Top- only when the uncertainty of in-distribution stimuli can be quantified by a smaller set. Despite earning higher bonuses, there is some evidence that participants who used prediction sets perceived lower worth compared to their Top- counterparts, as captured by their willingness-to-pay for the prediction displays. Therefore, uncertainty quantification does not guarantee more informed decision-making from users. We think designers of prediction displays should carefully consider the trade-off between set size and adaptiveness when using conformal prediction sets as a method to communicate prediction uncertainty for machine learning models in practical settings.
@@
Conformal prediction sets are more useful for hard OOD instances.
@@
One possible reason why prediction sets can exceed in performance for hard OOD stimuli is their adaptiveness. While the covariate shifts used in our experimental setting can severely affect the accuracy of Top- predictions, adaptive prediction sets under certain shifts can retain coverage by significantly increasing the set size (e.g., (Gendler et al., 2022)). While, the high coverage of prediction sets for OOD instances does not necessarily predict high accuracy of the AI-advised human, as the results for hard instances in Figure 6 demonstrate, even large set sizes for OOD instances provided some value to participants. Even when participants did not select the correct label from the set, when given a larger prediction set with hard OOD stimuli, participants tended to select labels that were closer to the correct one as measured by the shortest path length in the label space hierarchy. Hence, our findings highlight the value of conformal prediction sets in enhancing the reliability of predictions by machine learning models, especially when encountering OOD instances in real-world applications. Prediction display designers should leverage these sets’ adaptiveness and distribution-free properties.
Withholding predictions of an uncalibrated model can improve decision quality
@@
Consistent with prior work on AI-advised decision-making (e.g., (Green and Chen, 2019; Beede et al., 2020; Carton et al., 2020; Liu et al., 2021)), our results suggest that when a model is well-calibrated and more accurate than humans making judgments alone, users with access to its predictions can perform better than they would without the model, but not as well as the model alone for easier instances. When the model is poorly calibrated, we find that what sort of prediction display is shown affects whether people can do better by accessing model predictions. For instance, when given access to Top- displays with low coverage of the true label, users sometimes did worse than they would have by ignoring the display. We observed qualitative feedback such as “I used the Top- when it was impossible to find what the image was based on my own perception” and “in cases where I really was unsure and did not know what the picture was, I relied on the AI and went with their label”. The quality of decision-making when users are presented with a constrained set of label options is highly correlated with the magnitude of model calibration and hence performance can vary with the calibration of the uncertainty quantification. Only those who used RAPS reliably met or exceeded the expected performance without any prediction display. Our results underscore the pivotal role calibration plays and the potential value of treating the detection of instance type as a model’s learning problem (e.g., (Kaur et al., 2023)).
@@ 5.1. Limitations and Future Work
It is unclear whether a larger prediction set showing more possible labels can still be effective if the coverage is reduced. A natural follow-up study is to evaluate the efficacy of prediction sets in an OOD setting where coverage is disrupted at varying magnitudes. The adaptiveness of prediction sets (Romano et al., 2020; Angelopoulos et al., 2020) enhances a model’s resilience to random and common perturbations that can occur in real-world scenarios, such as image corruptions. However, future work may explore the use of intentional adversarial methods that can attack softmax (e.g., FGSM (Goodfellow et al., 2015) or PGD (Madry et al., 2017)). Combined with the finding from Straitouri and Rodriguez (2023) that constrained choices can lead to better performance based on sets derived from a calibrated model with a smaller size, we might see an opposite effect than what we observe, where people are less likely to pick a choice from a prediction set that has a lower chance of containing the true label.
We think willingness-to-pay is an imperfect comparator because our experimental design does not observe the performance of individual participants without a prediction display. Future work can further explore how users’ perceptions of the value of prediction displays and compare performance and perceived value within-subject. Additionally, it is unclear whether the coverage guarantee of conformal prediction sets is prone to misconceptions common to conventional confidence intervals, such as misinterpreting the coverage rate and disregarding labels not included within the set.
@@ 6. Conclusion
Uncertainty quantification for deep Neural Networks (NNs) has been challenging due to their lack of interpretability. We study the conformal prediction set as an alternative solution to quantify the prediction uncertainty of NNs and whether communicating prediction uncertainty can help human decision-making in AI-advised image labeling. We contribute the results of a large pre-registered online experiment, in which we evaluate the utility of conformal prediction sets against status quo Top- predictions using stimuli that are varied by both in- and out-of-distribution (OOD), difficulty, and size. Our work shows that when the model is well-calibrated, a smaller prediction set is most beneficial, as larger sets can overwhelm users and decrease performance due to increased cognitive load. However, when the model faces unexpected OOD instances, large prediction sets can be more useful than their Top- counterparts. Our work sheds light on how conformal prediction sets can be used to rigorously quantify uncertainty for machine learning models in practical applications and outlines potential avenues for future research.
@@ Appendix A Further Methodological Details
Corruption acc@1 acc@10 Coverage None 0.82 0.98 0.96 brightness 0.77 0.96 0.95 contrast 0.76 0.96 0.95 fog 0.72 0.93 0.94 jpeg compression 0.7 0.93 0.94 pixelate 0.66 0.9 0.92 motion blur 0.57 0.83 0.88 impulse noise 0.57 0.84 0.89 shot noise 0.56 0.83 0.89 gaussian noise 0.55 0.83 0.89 elastic transform 0.52 0.78 0.84 defocus blur 0.52 0.81 0.91 snow 0.51 0.79 0.86 frost 0.43 0.7 0.76 zoom blur 0.38 0.66 0.73 glass blur 0.29 0.55 0.72
Distribution of in-distribution cross-entropy loss with thresholds shown as vertical line
Corruption Size Difficulty Size1 Count SizeN Count Size Ratio SizeN Keep Size1 Include None small hard 539 2794 0.19 749 144 None small easy 13428 2962 4.53 1778 8060 defocus blur small hard 331 7750 0.04 870 37 defocus blur small easy 4571 2431 1.88 385 724 snow small hard 664 7319 0.09 827 75 snow small easy 4964 2554 1.94 451 877 frost small hard 993 9073 0.11 996 109 frost small easy 2896 1590 1.82 249 454 zoom blur small hard 722 9013 0.08 1066 85 zoom blur small easy 2925 1914 1.53 235 359 glass blur small hard 516 9973 0.05 1108 57 glass blur small easy 2025 1360 1.49 331 493
| Category | Open Codes | Description | ||
@@
| 1 | UNINFORMATIVE | Content that doesn’t provide any significant or relevant information | ||
@@
| 2 | HUMAN-AI COMPARISON | Participant analyzes differences and similarities between own’s decisions and AI predictions | ||
@@
| 3 | HUMAN KNOWLEDGE | Participant makes decisions based solely on own intuition, wisdom, or experience | ||
@@
| 4 | HUMAN KNOWLEDGE 1ST & AI ADVICE 2ND | Participant makes decision based on own’s intuition and experience first, supplemented by AI advice or suggestions | ||
@@
| 5 | AI ADVICE 1ST & HUMAN CONFIRM 2ND | Participant seeks advice from AI first and then confirms with own’s knowledge | ||
@@
| 6 | AI ADVICE WHEN HARD | Participant seeks advice from AI only and specifically in challenging or complex scenarios (e.g., blurry images, etc.) | ||
| 7 | TRUST AI PREDICTION(S) | Participant has confidence in the prediction made by AI | ||
| 8 | TRUST AI (RANKING OR CONFIDENCE) |
@@
| 9 | TOP PREDICTION FALLACY | Participant expects to see the true answer in top predictions, but they believe the top prediction is wrong | ||
@@
| 10 | DISAPPOINTED AT AI | Participant specifically mentions dissatisfaction with AI’s performance and does not trust AI’s prediction | ||
| 11 | SEARCH | Participant performs a general search without specifying | ||
@@
| 12 | SEARCH (BY DROPDOWN) | Participant types keywords in the input field and refers to the dropdown for matches | ||
@@
| 13 | SEARCH (BY KEYWORD) | Participant clicks “Search” to view a hierarchy tree that highlights matching categories | ||
| 14 | SEARCH (BY AI LABEL) | Participant clicks an AI-predicted label for a bottom-up search | ||
@@
| 15 | SEARCH (HIERARCHY NETWORK) | Participant expands or subtracts nodes/categories and navigates through the provided hierarchy network | ||
@@
| 16 | ELIMINATION | Participant tries to narrow down their options (to search in) by elimination | ||
@@
| 17 | REPRESENTATIVE IMAGES | Participant compares the task image against the given label-representative images | ||

### v6

@@ 4.4. Willingness-to-Pay
We present the distributions of elicited willingness-to-pay (WTP) measures in Figure 11 with summary statistics in Table 10 of Appendix B.3. Elicited WTPs across conditions are quite similar, suggesting that the measure may be noisy. Participants are willing to pay slightly more for Top-10 (; CI: ) than RAPS (; CI: ) and Top-1 (; CI: ).
Nonetheless, as detailed in Section 3.6, we quantify participants’ perceived utility for each prediction display using a measure we devised, called “willingness-to-overpay”. This proxy focuses on economic valuation, offering direct monetary comparisons between the monetary value of a prediction display that participants perceive (i.e., WTP) against how much additional expected bonus participants gained from accessing each prediction display.
| Condition | Accuracy |
| baseline | 0.41 | 1.64 | 0.00 | ||||
| Top-1 | 0.63 | 2.52 | 0.86 | ||||
| Top-10 | 0.64 | 2.56 | 0.88 | ||||
@@
This figure presents the bootstrapped 95% confidence intervals for the willingness-to-overpay measure by treatment conditions: Top-1, Top-10, and RAPS.
We note that this measure is imperfect as a comparator for any individual WTP because we cannot observe the counterfactual performance without a prediction display for each participant individually. In aggregate, however, Figure 8 suggests that participants may over-value the prediction displays to a similar extent across conditions. If anything, despite benefiting more from using the prediction sets, we see weak evidence that RAPS participants perceive a lower monetary value from the prediction set relative to Top- predictions. RAPS participants are willing to overpay by on average (CI ), slightly less than Top-1 participants (, CI ). Participants find Top-10 prediction most useful on average (; CI ).
@@ 4.5. Qualitative Analysis of Strategies
From our open coding, we find that participants employed different strategies between prediction displays (i.e., Figure 10 in Appendix B.1). In general, most participants reported consulting predictions when decision-making ( out of 450). Some participants reported first inspecting predictions and assessing their accuracy using their judgment (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150), while others report relying first on their own intuitions to form a general idea, and then seeing if it aligned with any predictions (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150). Very few participants chose to rely completely on their own intuitions because they find predictions inaccurate or not useful (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150). Others reported they would only consult predictions when the image was too hard to recognize and they had no clue (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150). Among participants who reported that they trusted AI prediction(s), some of them in Top- conditions placed higher trust on predictions with high softmax pseudo-probability (e.g., ) (Top-1: out of 23; Top-10: out of 35), while those in RAPS condition focused more on the top suggestions in the prediction set ( out of 26).
Collectively participants used all features of our interface, reporting that the dropdown ( out of 599), bottom-up ( out of 599), and keyword search ( out of 599) methods were useful in helping them identify a label, and the representative images were informative for aiding understanding of the meaning of labels ( out of 599). Of participants who viewed prediction displays, those in the Top-1 condition appeared most likely to use the label space hierarchy network to find a better match (60% out of 150, compared to 23% out of 150 and 28% out of 150).
@@ 5. Discussion
Conformal prediction has gained recent visibility as a distribution-free approach for quantifying uncertainty in model predictions. Our results suggest that decision-makers may be able to use prediction sets to their advantage in an AI-advised labeling task, but that their advantages over status quo Top- presentations vary with the properties of the setting. When the model has high accuracy and the test instances are in-distribution, we find that the size of the prediction set is a critical determinant of its utility in our setting. Smaller sets lead to performance on par with or slightly better than Top- displays while larger ones lead to slightly worse performance. In contrast, when the model encounters unforeseen and challenging OOD instances that can cause Top- predictions and their associated uncertainty scores to be unreliable (and which are also challenging for the humans), prediction sets offer some comparative advantage OOD, regardless of their set size, at least in a setting like ours where coverage guarantees are maintained. However, these advantages are relatively small, and may not always outweigh the benefits of instead using the calibration set instances in model training to further improve the classifier. We elucidate our main findings and discuss the broader implications of our study.
@@
Our results suggest that the utility of conformal prediction sets is linked to set size, presumably due to an underlying factor of cognitive load. A large set that embodies more uncertainty requires users to navigate and evaluate many incorrect predictions before finding the correct answer. This increase in cognitive load generally makes prediction sets less effective and may encourage more satisficing (Simon, 1956), resulting in a lower response accuracy. More rigorous uncertainty quantification does not always lead to better decision-making. Designers of prediction displays should carefully consider the trade-off between set size and adaptiveness when using conformal prediction sets to communicate prediction uncertainty for machine learning models in practical settings.
@@
Conformal prediction sets can be more useful for some hard OOD instances.
@@
One possible reason prediction sets can exceed in performance for hard OOD stimuli is their adaptiveness. While the covariate shifts used in our experimental setting can severely affect the accuracy of Top- predictions, adaptive prediction sets under certain shifts can retain coverage by significantly increasing the set size (e.g., (Gendler et al., 2022)). While the high coverage of prediction sets for OOD instances does not necessarily equate to the high accuracy of AI-advised humans, as the results for hard instances in Figure 6 demonstrate, even large set sizes for OOD instances provided some value to participants. Even if participants did not select the correct label from the set, when given larger prediction sets for hard OOD stimuli, they tended to select labels closer to the correct one, as measured by the shortest path length in the label space hierarchy. It is possible that the additional information about relative uncertainty provided by set size played a role in this advantage.
Withholding predictions of an uncalibrated model may improve decision quality
@@
Consistent with prior work on AI-advised decision-making (e.g., (Green and Chen, 2019; Beede et al., 2020; Carton et al., 2020; Liu et al., 2021)), our results suggest that when a model is well-calibrated and more accurate than humans alone, users with access to its predictions can perform better than without the model, but not as well as the model alone for easier instances. When the model is poorly calibrated, display type affects whether people can perform better by accessing the model predictions. For instance, when the Top- displays achieved lower coverage of the true label than the conformal guarantee, users sometimes performed worse than they would have done by ignoring the display. We observe qualitative feedback, such as “I used the Top- when it was impossible to find what the image was based on my own perception” and “In cases where I really was unsure and did not know what the picture was, I relied on the AI and went with their label.” Decision-making quality, when presented with a constrained set of choices, is correlated with the calibration of prediction information. Only those using RAPS matched the expected performance of having no prediction display in hard OOD instances with larger set sizes. Our results underscore the potential value of treating the detection of instance type as a model’s learning problem (e.g., (Kaur et al., 2023)).
@@ 5.1. Limitations and Future Work
It is unclear whether a larger prediction set showing more possible labels can still be effective if the coverage is reduced. A natural follow-up study is to evaluate the efficacy of prediction sets in an OOD setting where coverage is disrupted at varying magnitudes, as well as one where the human retains their relative judgment accuracy and only the AI accuracy is affected. The adaptiveness of the prediction sets (Romano et al., 2020; Angelopoulos et al., 2020) enhances a model’s resilience to random and common perturbations that can occur in real-world scenarios, such as image corruptions. However, future work may explore the use of intentional adversarial methods that can attack softmax (e.g., FGSM (Goodfellow et al., 2015) or PGD (Madry et al., 2017)). Combined with the finding from Straitouri and Rodriguez (2023) that constrained choices can lead to better performance based on smaller sets derived from a calibrated model, we might see an opposite effect to what we observed, where people are less likely to pick a choice from a prediction set that has a lower chance of containing the true label. While our experiment isolated the effect of the conformal prediction set and guarantee without performing other interpretability manipulations to the Top- conditions, future work might also consider comparing conformal prediction sets to post-hoc calibrated softmax with Top- displays.
Our experimental design, which does not observe individual participant performance without a prediction display, makes willingness-to-pay an imperfect comparator. Future work should delve into how users perceive the value of prediction displays and conduct within-subject comparisons of both performance and perceived value. Additionally, it is unclear whether the conformal coverage guarantee is prone to misconceptions common to conventional confidence intervals, such as misinterpreting the coverage rate and disregarding labels not included within the set.
@@ 6. Conclusion
Uncertainty quantification for deep neural networks (NNs) has been challenging due to their black-box nature. We evaluated whether communicating uncertainty via conformal prediction sets can improve human decision-making in AI-advised image labeling. We contribute the results of a large online experiment, in which we evaluate the utility of conformal prediction sets against status quo Top- predictions across a diverse range of stimuli, varied by in-distribution or OOD, level of difficulty, and set size. Our work shows that when the model is well-calibrated, a smaller prediction set is most beneficial, as larger sets can overwhelm users and decrease performance due to increased cognitive load. However, when the model faces unexpected OOD instances that are also more difficult for humans, larger prediction sets can be more useful than their Top- counterparts, at least when coverage remains high. Our work sheds light on how conformal prediction sets can be used to rigorously quantify uncertainty for machine learning models in practical applications and outlines potential avenues for future research.
@@ Appendix A Further Methodological Details
We present the distribution of the cross-entropy loss for stimuli that are in-distribution in Figure 9. Table 6 shows WRN’s Top- prediction accuracy and conformal coverage for in-distribution stimuli and 15 types of covariate shifts. Table 8 outlines the sampling correction for images with a deterministic set, while Table 9 explains the open codes used to categorize participants’ strategies. Furthermore, Table 7 compares the performance of the Human-AI team, humans working alone, and AI standalone accuracy.
This figure illustrates the distribution of cross-entropy loss for in-distribution predictions. To facilitate comparison of the median cross-entropy values, we use two visual cues: an orange dotted line represents the median when including images with a set size of one, while a red solid line denotes the median after excluding these images.
| Corruption | Coverage | ||
| Top-1 | Top-10 | RAPS | |
| none | 0.82 | 0.98 | 0.96 |
| brightness | 0.77 | 0.96 | 0.95 |
| contrast | 0.76 | 0.96 | 0.95 |
| fog | 0.72 | 0.93 | 0.94 |
| jpeg compression | 0.7 | 0.93 | 0.94 |
| pixelate | 0.66 | 0.9 | 0.92 |
| motion blur | 0.57 | 0.83 | 0.88 |
| impulse noise | 0.57 | 0.84 | 0.89 |
| shot noise | 0.56 | 0.83 | 0.89 |
| gaussian noise | 0.55 | 0.83 | 0.89 |
| elastic transform | 0.52 | 0.78 | 0.84 |
| defocus blur | 0.52 | 0.81 | 0.91 |
| snow | 0.51 | 0.79 | 0.86 |
| frost | 0.43 | 0.7 | 0.76 |
| zoom blur | 0.38 | 0.66 | 0.73 |
| glass blur | 0.29 | 0.55 | 0.72 |
| Condition | Covariate Shift | Accuracy | ||
| Human-AI Team | AI Baseline | Human Baseline | ||
| Top-1 | in | 0.74 | 0.73 | 0.53 |
| out | 0.59 | 0.52 | 0.38 | |
| Top-10 | in | 0.76 | 0.73 | 0.53 |
| out | 0.59 | 0.52 | 0.38 | |
| RAPS | in | 0.73 | 0.73 | 0.53 |
| out | 0.63 | 0.52 | 0.38 | |
| Corruption | Difficulty | Size1 Count | SizeN Count | Size Ratio | SizeN Keep | Size1 Included |
| none | hard | 539 | 2794 | 0.19 | 749 | 144 |
| none | easy | 13428 | 2962 | 4.53 | 1778 | 8060 |
| defocus blur | hard | 331 | 7750 | 0.04 | 870 | 37 |
| defocus blur | easy | 4571 | 2431 | 1.88 | 385 | 724 |
| snow | hard | 664 | 7319 | 0.09 | 827 | 75 |
| snow | easy | 4964 | 2554 | 1.94 | 451 | 877 |
| frost | hard | 993 | 9073 | 0.11 | 996 | 109 |
| frost | easy | 2896 | 1590 | 1.82 | 249 | 454 |
| zoom blur | hard | 722 | 9013 | 0.08 | 1066 | 85 |
| zoom blur | easy | 2925 | 1914 | 1.53 | 235 | 359 |
| glass blur | hard | 516 | 9973 | 0.05 | 1108 | 57 |
| glass blur | easy | 2025 | 1360 | 1.49 | 331 | 493 |
| Category | Open Codes | Description |
@@
| 1 | UNINFORMATIVE | Content that doesn’t provide any significant or relevant information |
@@
| 2 | HUMAN-AI COMPARISON | Participant analyzes differences and similarities between own’s decisions and AI predictions |
@@
| 3 | HUMAN KNOWLEDGE | Participant makes decisions based solely on own intuition, wisdom, or experience |
@@
| 4 | HUMAN KNOWLEDGE 1ST & AI ADVICE 2ND | Participant makes decision based on own’s intuition and experience first, supplemented by AI advice or suggestions |
@@
| 5 | AI ADVICE 1ST & HUMAN CONFIRM 2ND | Participant seeks advice from AI first and then confirms with own’s knowledge |
@@
| 6 | AI ADVICE WHEN HARD | Participant seeks advice from AI only and specifically in challenging or complex scenarios (e.g., blurry images, etc.) |
| 7 | TRUST AI PREDICTION(S) | Participant has confidence in the prediction made by AI |
| 8 | TRUST AI (RANKING OR CONFIDENCE) | Participant relies on AI’s ranked suggestions or its confidence level in predictions. Participant mentions looking only at top predictions in RAPS and Top-10, while in Top-1 referring to specific confidence percentages |
@@
| 9 | TOP PREDICTION FALLACY | Participant expects to see the true answer in top predictions, but they believe the top prediction is wrong |
@@
| 10 | DISAPPOINTED AT AI | Participant specifically mentions dissatisfaction with AI’s performance and does not trust AI’s prediction |
| 11 | SEARCH | Participant performs a general search without specifying |
@@
| 12 | SEARCH (BY DROPDOWN) | Participant types keywords in the input field and refers to the dropdown for matches |
@@
| 13 | SEARCH (BY KEYWORD) | Participant clicks “Search” to view a hierarchy tree that highlights matching categories |
| 14 | SEARCH (BY AI LABEL) | Participant clicks an AI-predicted label for a bottom-up search |
@@
| 15 | SEARCH (HIERARCHY NETWORK) | Participant expands or subtracts nodes/categories and navigates through the provided hierarchy network |
@@
| 16 | ELIMINATION | Participant tries to narrow down their options (to search in) by elimination |
@@
| 17 | REPRESENTATIVE IMAGES | Participant compares the task image against the given label-representative images |